# FlowR: Flowing from Sparse to Dense 3D Reconstructions

### v1

@@ Abstract
3D Gaussian splatting enables high-quality novel view synthesis (NVS) at real-time frame rates. However, its quality drops sharply as we depart from the training views. Thus, dense captures are needed to match the high-quality expectations of some applications, e.g. Virtual Reality (VR). However, such dense captures are very laborious and expensive to obtain. Existing works have explored using 2D generative models to alleviate this requirement by distillation or generating additional training views. These methods are often conditioned only on a handful of reference input views and thus do not fully exploit the available 3D information, leading to inconsistent generation results and reconstruction artifacts. To tackle this problem, we propose a multi-view, flow matching model that learns a flow to connect novel view renderings from possibly sparse reconstructions to renderings that we expect from dense reconstructions. This enables augmenting scene captures with novel, generated views to improve reconstruction quality. Our model is trained on a novel dataset of 3.6M image pairs and can process up to 45 views at $540\times 960$ resolution (91K tokens) on one H100 GPU in a single forward pass. Our pipeline consistently improves NVS in sparse- and dense-view scenarios, leading to higher-quality reconstructions than prior works across multiple, widely-used NVS benchmarks.
@@ 1 Introduction
3D reconstruction is the process of estimating the geometry and appearance of a 3D scene from a set of 2D images of that scene. Given a large and dense enough set of images of the 3D scene, modern 3D reconstruction methods such as neural radiance fields (NeRF) [36] and 3D Gaussian splatting (3DGS) [24] can reconstruct 3D scenes in such high quality that the resulting 3D representations can be rendered into novel views that are almost indistinguishable from reality. This is the task of novel view synthesis (NVS), and it enables various applications such as immersive VR experiences inside of photo-realistically captured real scenes.
However, for these methods to achieve such results, an extremely large number of captured images is needed for a given 3D scene. Capturing such a large number of images is laborious and not always possible. Thus, one of the core challenges in 3D reconstruction and novel view synthesis is how to build an algorithm that can achieve equivalently good results with far fewer images, while also being able to take advantage of a large set of images when they are available.
In this paper, we present our FlowR method to tackle this problem. FlowR consists of two parts: (a) a robust initial reconstruction pipeline based on 3DGS [24] but designed for both sparse- and dense-view settings. We use MASt3R [27] to estimate tracked correspondences which are then used to triangulate an initial point cloud; and (b) a data densification procedure, which uses flow matching [32] to generate high-quality extra views which can be used to improve reconstruction.
Although several previous approaches  [43, 56, 58, 19, 64, 12, 33] have used generative models to generate novel views to improve reconstruction, we are, to the best of our knowledge, the first to pose the problem as a flow matching problem, learning the transformation between the distribution of (incorrect) rendered images and ground truth images. Flow matching is a paradigm for generative modeling where the model learns a velocity field that can be used to map samples from a noise distribution to samples from the data distribution. However, in this work, instead of modeling the velocity field between noise and data, we model the velocity field between the incorrect novel view renderings and the respective real images of that viewpoint (see Fig. 1). In this way, if we already have enough dense input images that our initial reconstruction is good enough for a particular view, the flow matching model can simply learn not to change the input. This formulation ensures that the generative model does not generate unnecessary new details that are inconsistent and conflict with existing scene content, and thus results in sharp scene details, avoiding blurred-out averages of inconsistent generations.
@@
  • We develop a scalable and robust 3D reconstruction pipeline that produces semi-dense, metric-scale 3D reconstructions from arbitrary view distributions and collect a large-scale dataset with 3.6M rendered and ground-truth image pairs.
  • We propose a flow matching formulation that directly incorporates novel-view renderings from initial reconstructions as a surrogate for the initial “noise” distribution.
  • We demonstrate that our trained model enhances the quality of 3D reconstructions in both sparse and dense view scenarios by simply incorporating novel-view renderings fixed by our flow matching model into the reconstruction process.
We demonstrate that our approach leads to superior novel view synthesis results across various benchmarks, outperforming prior methods in sparse- and dense-view scenarios.
@@ 2 Preliminaries
3D Gaussian splatting [24] generalizes point clouds as a representation of 3D scenes.
Representation. Let $\mathcal{I}\coloneqq\{\mathcal{I}_{i}\coloneqq(\mathtt{I}_{i},\mathbf{P}_{i},\mathbf{K}_{i})\}_{i=1}^{N}$ denote a set of $N$ input images $\mathtt{I}_{i}$ with camera to world transformation $\mathbf{P}_{i}\coloneqq[\mathbf{R}_{i}|\mathbf{t}_{i}]$ and intrinisc matrix $\mathbf{K}_{i}$. We are interested in representing the underlying 3D scene as a set of 3D Gaussian primitives $\mathcal{G}\coloneqq\{\mathfrak{g}_{k}\}_{k=1}^{K}$. Each 3D Gaussian $\mathfrak{g}_{k}\in\mathcal{G}$ is parametrized by $\mathfrak{g}_{k}\coloneqq\{\boldsymbol{\mu}_{k},\mathbf{s}_{k},\mathbf{q}_{k},\alpha_{k},\mathbf{c}_{k}\}$, i.e. position, scale, rotation, opacity and view-dependent color, respectively. The Gaussian kernel takes the following form
@@
$\mathbf{c}(\mathbf{p}^{\prime})\coloneqq\sum_{k=0}^{K}\mathbf{c}_{k}w_{k}\prod_{j=0}^{k-1}(1-w_{j})\,,$ (2)
where $\quad w_{k}\coloneqq\alpha_{k}\mathfrak{g}^{c}_{k}(\mathbf{p}^{\prime})$.
@@
where $\hat{\mathtt{I}}$ is a rendered image as per Eq. 2 from the same viewpoint of training image $\mathtt{I}$, $\lambda_{\text{ssim}}\coloneqq 0.2$, and $\operatorname{SSIM}(\cdot)$ is the structural similarity loss [55]. The optimization is interleaved with an adaptive density control (ADC) mechanism [24] that prunes transparent primitives, and splits and clones primitives in over- and under-reconstructed areas, respectively.
@@ 3 Method
Our goal is to improve the quality of an initial reconstruction of a scene created from a possibly sparse set of posed images. We exploit flow matching to improve the quality of novel view renderings from the initial reconstruction, which serve as auxiliary training data to fit an improved 3D representation. Although our method works potentially with different scene representations, it is convenient to adopt Gaussian primitives because they can be trained fast. In fact, since our approach is data-driven, we need to construct a dataset of reconstructions under different sparsity levels, and Gaussian splatting ensures better scalability for this purpose. In Sec. 3.1, we describe our robust 3D reconstruction method that serves as a basis for generating our dataset of 3D reconstructions and provides the initial scene reconstructions. Sec. 3.2 describes how we employ flow matching to improve novel-view renderings from the initial reconstructions, which, in turn, are used to improve our reconstruction results in Sec. 3.3. See Fig. 2 for an illustration.
@@ 3.1 Robust 3D Reconstruction
We aim to devise a semi-dense, metric-scale 3D reconstruction pipeline that is robust to the input view distribution, i.e. sparse and dense view distributions ranging from a few to thousands of input views. To this end, we combine recent advances in learning-based structure-from-motion (SfM) and monocular depth estimation with classical SfM tools.
First, we define a sparse co-visibility graph $G_{\text{vis}}\coloneqq(\mathcal{I},\mathcal{E})$ on the set of posed input images $\mathcal{I}$ similar to [14]. We sample $\sqrt{N}$ keyframes using farthest point sampling. The edges $\mathcal{E}\subset\mathcal{I}\times\mathcal{I}$ are constructed by densely connecting all keyframes and connecting other nodes to their closest keyframe and to the $k$ nearest neighbors. For both keyframes selection and edge construction we consider the simple image co-visibility metric, $\mathcal{S}(\mathcal{I}_{i},\mathcal{I}_{j})\coloneqq 1-\frac{\mathcal{D}(\mathcal{I}_{i},\mathcal{I}_{j})}{\max_{ij}\mathcal{D}(\mathcal{I}_{i},\mathcal{I}_{j})}$, where
@@
Fitting 3D Gaussians. Using this reconstruction as an initialization, we fit 3D Gaussians $\mathcal{G}$ from the input images with a short schedule. We train for $5$k steps with a warm-up phase of 200 steps. After warm-up, we use ADC with AbsGrad [61] metric until step $2.5$k. Compared to previous works [17] that use 3D points obtained from global pointcloud alignment [27] as initialization, our pipeline has two advantages. First, it refines the geometry by using track constraints. Second, it removes duplicated and inconsistent 3D points from the initialization, improving performance and scalability (see Sec. 4.2).
Dataset generation. We leverage our robust reconstruction pipeline and recent large-scale NVS benchmarks [29, 62] to generate a dataset of rendered and ground-truth image pairs. For each scene, we first select an input view subset with a varying degree of view sparsity to achieve a large variety of view distributions (see Sec. A.1). Next, we utilize all views not in this subset as privileged information for our flow matching model. We reconstruct the scenes from the input view subset with our aforementioned pipeline and create rendering, ground-truth pairs from all other available views, producing a training dataset of $3.6$M image pairs from $10.3$k sequences. Processing all scenes took about $1.1$k GPU hours or 6.5 minutes on average per scene on budget GPUs like RTX 2080 Ti.
@@ 3.2 Flowing from Sparse to Dense Reconstructions
Velocity field $\boldsymbol{v}_{\theta}(\mathbf{z},\mathbf{y},t)$. We implement the velocity field of the flow matching model using the diffusion transformer (DiT) [15, 40] architecture. All image latents provided as input, namely $\mathbf{z}$ and the encoded source images from $\mathbf{y}_{\text{src}}$, are first split into $2\times 2$ patches. We project each patch to the token dimension and add a 2D positional encoding. In addition, we encode the index $i$ of the image from which each token originates. To this end, we adopt a 1D sinusoidal encoding $\gamma(i)$ [52]. During training, the view index is chosen randomly from $[0,1000]$, whereas at inference the view indices are $[0,N]$ where $N$ the number of images for a single forward pass. Originally from LLM literature [10], this technique enables inference with very long context lengths and has recently been applied to 3D reconstruction [59]. The sequence of tokens is then processed with a series of transformer blocks to predict the output velocity tensor. Transformer blocks consist of a self-attention and a feed-forward residual layer. To jointly process multiple reference and target views while leveraging large-scale text-to-image pre-training, we adapt the DiT block as follows. We first collapse the view dimension into the batch dimension, keeping the self-attention layer equivalent to the pre-trained image model. Then, we collapse the view dimension into the spatial dimension. We add the image index encoding $\gamma(i)$ to the state, concatenate it with the ray map embeddings, and linearly project the concatenated tokens back to the original dimension. We then insert a second self-attention layer to enable full multi-view attention. Finally, before adding the output to the prior state of the block, we insert a zero linear layer [67] to keep the pre-trained initialization intact. See Sec. A.2 for an illustration.
@@ 3.3 Improving the 3D Reconstruction
Generation of target views $\mathcal{P}_{\text{tgt}}$. We first select a set of camera poses suitable for augmenting the initial input views. We distinguish between scenes where sparse input views lie on a continuous trajectory, i.e. input views are ordered, and unordered photo collections. If input views are ordered, we fit a smooth trajectory to the source camera poses using B-spline basis functions. If not, we use the distance metric in Eq. 8 to select reference poses with farthest point sampling. For each reference pose, we generate candidate points on a sphere with a random radius using Fibonacci sphere sampling [48]. We select the point with the maximum distance to the source cameras and apply a small perturbation to the reference view orientation to obtain the target camera pose. Finally, we filter target camera poses with too many points close to the camera or too few points inside the view frustum using the initial 3D reconstruction results, yielding the final set of target poses $\mathcal{P}_{\text{tgt}}$.
@@
Refined reconstruction $\mathcal{G}_{\text{ref}}$. Finally, we fit an improved 3D scene representation $\mathcal{G}_{\text{ref}}$ on the union $\mathcal{I}_{\text{src}}\cup\mathcal{I}_{\text{tgt}}$ of source and target novel-view posed images. However, the two sets of images undergo different loss functions. Namely, we apply the loss function in Eq. 3 to images in $\mathcal{I}_{\text{src}}$ and the following loss function to images in $\mathcal{I}_{\text{tgt}}$:
@@ 4 Experiments
Experimental setup. We compare to prior art across multiple benchmarks, i.e. DL3DV140 [29], ScanNet+++++ + [62], and Nerfbusters [57]. For DL3DV140, we first split each sequence into training and evaluation views and create two training splits by sampling $k\in\{12,24\}$ equally spaced views from the training trajectory. We evaluate each sequence at the original $540\times 960$ resolution. For ScanNet+++++ +, we use the validation set, follow the official training and testing splits, and evaluate at $640\times 960$ resolution. For Nerfbusters, we follow the original evaluation protocol in [57]. We measure view quality in terms of both low-level similarity via PSNR and SSIM [55], and high-level, perceptual similarity using LPIPS [68].

### v2

@@ Abstract
3D Gaussian splatting enables high-quality novel view synthesis (NVS) at real-time frame rates. However, its quality drops sharply as we depart from the training views. Thus, dense captures are needed to match the high-quality expectations of applications like Virtual Reality (VR). However, such dense captures are very laborious and expensive to obtain. Existing works have explored using 2D generative models to alleviate this requirement by distillation or generating additional training views. These models typically rely on a noise-to-data generative process conditioned only on a handful of reference input views, leading to hallucinations, inconsistent generation results, and subsequent reconstruction artifacts. Instead, we propose a multi-view, flow matching model that learns a flow to directly connect novel view renderings from possibly sparse reconstructions to renderings that we expect from dense reconstructions. This enables augmenting scene captures with consistent, generated views to improve reconstruction quality. Our model is trained on a novel dataset of 3.6M image pairs and can process up to 45 views at $540\times 960$ resolution (91K tokens) on one H100 GPU in a single forward pass. Our pipeline consistently improves NVS in sparse- and dense-view scenarios, leading to higher-quality reconstructions than prior works across multiple, widely-used NVS benchmarks.
@@ 1 Introduction
3D reconstruction is the process of estimating the 3D scene geometry and appearance from a set of 2D images. Given a large and dense enough set of images, modern 3D reconstruction methods such as neural radiance fields (NeRF) [37] and 3D Gaussian splatting (3DGS) [25] produce 3D representations that can be rendered into novel views that are almost indistinguishable from reality. This is the task of novel view synthesis (NVS), and it enables various applications such as immersive VR experiences inside captured real-world scenes. However, for these methods to achieve such a high level of photorealism, they need an extremely large number of captured images per scene. Obtaining these images is laborious and not always possible. Thus, one of the core challenges in 3D reconstruction and novel view synthesis is how to build an algorithm that can achieve equally good results with far fewer images, while also being able to take advantage of a large set of images when they are available.
In this paper, we tackle this problem with our FlowR method. FlowR consists of two parts: (a) a robust initial reconstruction pipeline based on 3DGS [25] but designed for both sparse- and dense-view settings; and (b) a data densification procedure, which uses flow matching [33] to generate high-quality extra views that can be used to improve reconstruction.
Several previous approaches [45, 59, 61, 19, 67, 12, 34] have used generative models to generate novel views to improve reconstruction. These works rely on diffusion [21] or flow matching [31] models that map samples between a data and a noise distribution. As such, these works use a noise-to-data generative process, usually conditioned on initial input images or renderings (Fig. 1, top). We hypothesize that a generative process that directly maps samples from a conditional source distribution of rendered images to the target distribution of ground truth images is preferable to conditional noise-to-data generation. Additionally, we observe that flow matching is a paradigm for generative modeling where the model learns a velocity field that transports samples from any source to a target distribution. Therefore, in this work, we pose the problem as a flow matching problem where, instead of modeling the velocity field between noise and data, we model the velocity field between the distribution of incorrect novel view renderings and the respective real images of that viewpoint (Fig. 1, bottom).
In this way, if we already have enough dense input images that our initial reconstruction is good enough for a particular view, the flow matching model can simply learn not to change the input. This formulation ensures that the generative model does not hallucinate unnecessary new details that are inconsistent and conflict with existing scene content, and thus results in sharp scene details, avoiding blurred-out averages of inconsistent generations.
@@
  • We develop a fast initial 3D reconstruction pipeline that produces 3DGS representations from arbitrary view distributions and use it to collect a large-scale dataset of 3.6M rendered and ground-truth image pairs.
  • We propose a flow matching formulation that directly incorporates novel-view renderings from initial reconstructions as a surrogate for the initial “noise” distribution and use it to train a multi-view flow matching model that enhances novel-view renderings.
  • We demonstrate that our trained model enhances the quality of 3D reconstructions in both sparse and dense view scenarios by simply incorporating novel-view renderings enhanced by our flow matching model into the reconstruction process.
Our approach outperforms prior methods in both sparse- and dense-view scenarios across three NVS benchmarks.
@@ 2 Preliminaries
3D Gaussian splatting [25] describes a method for NVS from a set of calibrated images.
Representation. Let $\mathcal{I}\coloneqq\{\mathcal{I}_{i}\coloneqq(\mathtt{I}_{i},\mathbf{P}_{i},\mathbf{K}_{i})\}_{i=1}^{N}$ denote a set of $N$ input images $\mathtt{I}_{i}$ with camera-to-world transformation $\mathbf{P}_{i}\coloneqq[\mathbf{R}_{i}|\mathbf{t}_{i}]$ and intrinisc matrix $\mathbf{K}_{i}$. We are interested in representing the underlying 3D scene as a set of 3D Gaussian primitives $\mathcal{G}\coloneqq\{\mathfrak{g}_{k}\}_{k=1}^{K}$. Each 3D Gaussian $\mathfrak{g}_{k}\in\mathcal{G}$ is parametrized by $\mathfrak{g}_{k}\coloneqq\{\boldsymbol{\mu}_{k},\mathbf{s}_{k},\mathbf{q}_{k},\alpha_{k},\mathbf{c}_{k}\}$, i.e. position, scale, rotation, opacity and view-dependent color, respectively. The Gaussian kernel takes the following form
@@
$\mathbf{c}(\mathbf{p}^{\prime})\coloneqq\sum_{k=0}^{K}\mathbf{c}^{\prime}_{k}w_{k}\prod_{j=0}^{k-1}(1-w_{j})\,$, (2)
where $\quad w_{k}\coloneqq\alpha_{k}\mathfrak{g}^{c}_{k}(\mathbf{p}^{\prime})$ and $\mathbf{c}^{\prime}_{k}$ is the evaluated spherical harmonics function of $\mathbf{c}_{k}$ at the view direction of $c$.
@@
where $\hat{\mathtt{I}}$ is a rendering as per Eq. 2 from the viewpoint of training image $\mathtt{I}$, $\lambda_{\text{ssim}}\coloneqq 0.2$, and $\operatorname{SSIM}(\cdot)$ is the structural similarity loss [58]. The optimization is interleaved with an adaptive density control (ADC) mechanism [25] that prunes transparent primitives, and splits and clones primitives in over- and under-reconstructed areas, respectively.
@@ 3 Method
Our goal is to improve the quality of an initial reconstruction of a scene created from a possibly sparse set of posed images. We exploit flow matching to improve the quality of novel view renderings from the initial reconstruction, which serve as auxiliary training data to fit an improved 3D representation. Although our method could work with different scene representations, it is convenient to adopt Gaussian primitives because they can be trained fast. In fact, since our approach is data-driven, we need to construct a dataset of reconstructions under different sparsity levels, and Gaussian splatting ensures better scalability for this purpose. In Sec. 3.1, we describe our robust 3D reconstruction method that serves as a basis for generating our dataset of 3D reconstructions and provides the initial scene reconstructions. Sec. 3.2 describes how we employ flow matching to improve novel-view renderings from the initial reconstructions, which, in turn, are used to improve our reconstruction results in Sec. 3.3. See Fig. 2 for an illustration.
@@ 3.1 Robust 3D Reconstruction
We aim to devise a 3D reconstruction pipeline that produces semi-dense, metric-scale point clouds from arbitrary input view distributions, i.e. ranging from a few sparse to thousands of dense views. To this end, we combine recent advances in learning-based structure-from-motion (SfM) and monocular depth estimation with classical SfM tools.
First, we define a sparse co-visibility graph $G_{\text{vis}}\coloneqq(\mathcal{I},\mathcal{E})$ on the set of posed input images $\mathcal{I}$ similar to [14]. We use keyframing to limit the number of view pairs for matching and thus enable a scalable 3D reconstruction process. In particular, we sample $\sqrt{N}$ keyframes using farthest point sampling. The edges $\mathcal{E}\subset\mathcal{I}\times\mathcal{I}$ are constructed by densely connecting all keyframes and connecting other nodes to their closest keyframe and their $k$ nearest neighbors. For both keyframes selection and edge construction we consider the simple image co-visibility metric, $\mathcal{S}(\mathcal{I}_{i},\mathcal{I}_{j})\coloneqq 1-\frac{\mathcal{D}(\mathcal{I}_{i},\mathcal{I}_{j})}{\max_{ij}\mathcal{D}(\mathcal{I}_{i},\mathcal{I}_{j})}$, where
@@
Fitting 3D Gaussians. Using this reconstruction as an initialization, we fit 3D Gaussians $\mathcal{G}$ from the input images with a short schedule. We train for 55k steps with a warm-up phase of 200 steps. After warm-up, we use ADC with AbsGrad [64] metric until step 2.52.5k. Compared to previous works [17] that use 3D points obtained from global pointcloud alignment [28] as initialization, our pipeline has two advantages. First, it refines the geometry by using track constraints. Second, it removes duplicated and inconsistent 3D points from the initialization, improving performance and scalability (see Sec. 4.2).
Dataset generation. We leverage our robust reconstruction pipeline and two recent large-scale NVS benchmarks [30, 65] to generate a dataset of rendered and ground-truth image pairs. For each scene, we first select an input view subset with a varying degree of view sparsity to achieve a large variety of view distributions (see Sec. A.1). Next, we utilize all views not in this subset as privileged information for our flow matching model. We reconstruct the scenes from the input view subset with our aforementioned pipeline and create rendering, ground-truth pairs from all other available views, producing a training dataset of 3.63.6M image pairs from 10.310.3k sequences. Processing all scenes took about 1.11.1k GPU hours or 6.5 minutes on average per scene on budget GPUs like RTX 2080 Ti.
@@ 3.2 Flowing from Sparse to Dense Reconstructions
Velocity field $\boldsymbol{v}_{\theta}(\mathbf{z},\mathbf{y},t)$. We input the target and source views as a single set of tokens to the flow matching model implemented as a diffusion transformer (DiT) [15, 42]. All image latents provided as input, namely $\mathbf{z}$ and the encoded source images from $\mathbf{y}_{\text{src}}$, are first split into $2\times 2$ patches. We project each patch to the token dimension and add a 2D positional encoding. In addition, we encode the index $i$ of the image from which each token originates. To this end, we adopt a 1D sinusoidal encoding $\gamma(i)$ [55]. During training, the view index is chosen randomly from $[0,1000]$, whereas at inference the view indices are $[0,N]$ where $N$ the number of images for a single forward pass. Originally from LLM literature [10], this technique enables inference with very long context lengths and has recently been applied to 3D reconstruction [62]. The sequence of tokens is then processed with a series of transformer blocks to predict the output velocity tensor. To share information across multiple reference and target views while leveraging large-scale text-to-image pre-training, we adapt the DiT block as follows. We first collapse the view dimension into the batch dimension, keeping the self-attention layer equivalent to the pre-trained image model. Then, we collapse the view dimension into the spatial dimension. We add the image index encoding $\gamma(i)$ to the state, concatenate it with the ray map embeddings, and linearly project the concatenated tokens back to the original dimension. We then insert a second self-attention layer to enable full multi-view attention. Finally, before adding the output to the prior state of the block, we insert a zero linear layer [70] to keep the pre-trained initialization intact. See Sec. A.2 for an illustration.
@@ 3.3 Improving the 3D Reconstruction
Generation of target views $\mathcal{P}_{\text{tgt}}$. We first select a set of camera poses suitable for augmenting the initial input views. We distinguish between scenes where sparse input views lie on a continuous trajectory, i.e. input views are ordered, and unordered photo collections. If input views are ordered, we model a smooth trajectory using B-spline basis functions from a sparse set of control points (source camera poses) and sample the target poses from it. If not, we use the distance metric in Eq. 8 to select reference poses with farthest point sampling. For each reference pose, we generate candidate points on a sphere with a random radius using Fibonacci sphere sampling [51]. We select the point with the maximum distance to the source cameras and apply a small perturbation to the reference view orientation to obtain the target camera pose. Finally, we filter target camera poses with too many points close to the camera or too few points inside the view frustum using the initial 3D reconstruction results, yielding the final set of target poses $\mathcal{P}_{\text{tgt}}$.
@@
Refined reconstruction $\mathcal{G}_{\text{ref}}$. We fit an improved 3D scene representation $\mathcal{G}_{\text{ref}}$ on the union $\mathcal{I}_{\text{src}}\cup\mathcal{I}_{\text{tgt}}$ of source and target novel-view posed images. However, the two sets of images undergo different loss functions because views in $\mathcal{I}_{\text{tgt}}$ could exhibit small imperfections, e.g., due to VAE compression artifacts. Namely, we apply Eq. 3 to images in $\mathcal{I}_{\text{src}}$ and the following loss function to images in $\mathcal{I}_{\text{tgt}}$:
@@ 4 Experiments
Experimental setup. We compare to prior art across multiple benchmarks, i.e. DL3DV140 [30], ScanNet++++ [65], and Nerfbusters [60]. For DL3DV140, we first split each sequence into training and evaluation views and create two training splits by sampling $k\in\{12,24\}$ equally spaced views from the training trajectory. We evaluate each sequence at the original $540\times 960$ resolution. For ScanNet++++, we use the validation set, follow the official training and testing splits, and evaluate at $640\times 960$ resolution. For Nerfbusters, we follow the original evaluation protocol in [60]. We measure view quality in terms of both low-level similarity via PSNR and SSIM [58], and high-level, perceptual similarity using LPIPS [71].